If (S,*) is a.semigroupand x € s show that (S,∆) is a semigroup if a∆b =a*x*b


1
Expert's answer
2021-09-29T08:49:48-0400

Let "(S,*)" be a semigroup and "x\in S". Let us show that "(S,\u2206)" is a semigroup if "a\u2206b =a*x*b."

Since "x\in S", then "a,b\in S" imply that "a*x*b\in S," and hence "a∆b\in S." It follows that the operation "\u2206" is defined on the set "S."

Taking into account that

"(a∆b)∆c=(a*x*b)∆c=(a*x*b)*x*c\\=a*x*(b*x*c)=a*x*(b∆c)=a∆(b∆c),"

we conclude that the operation "\u2206" is associative, and hence "(S,\u2206)" is a semigroup.